OpenStudy (anonymous):

MEDAL FOR THE BEST ANSWER The line plot shows the part of an hour that Tamika practiced the piano each day last week . How many hours did Tamika practice the piano last week? Hours practiced X X X X X X X ------------------ 0 1/4 1/2 3/4 1 A. 4 hours B. 1 1/2 hours C. 3 3/4 hours D. 4 1/2 hours

OpenStudy (calculusxy):

Are the x's located above 1/4, 1/2, and 3/4?

OpenStudy (anonymous):

yes

OpenStudy (calculusxy):

My best way: 1/4*1= 1/2*3= 3/4*3= Find the products, then add.

OpenStudy (calculusxy):

Understand?

OpenStudy (anonymous):

um sort of... all i got was 48/32 but they dont have that answer!

OpenStudy (calculusxy):

Okay I am checking.

OpenStudy (calculusxy):

What are the products of 1/4*1, 1/2*3, and 3/4*3?

OpenStudy (anonymous):

i got 8/32 + 16/32 + 24/32

OpenStudy (calculusxy):

Multiply the numbers in the numerator, then divide it by the number in the denominator.

OpenStudy (anonymous):

ok thnx!! :D

OpenStudy (calculusxy):

Then add.